# Transversal Hamilton cycles in digraph collections

### v1

@@ 2.1 Regularity for digraph collections
For every integer $L_{0}\geq 1$ and every $\epsilon,\delta>0$, there is an $n_{0}=n_{0}(\epsilon,\delta,L_{0})$ such that for every $d\in[0,1)$ and every digraph collection $\mathcal{D}=\{D_{i}:i\in\mathcal{C}\}$ on vertex set $V$ of size $n\geq n_{0}$ with $\delta n\leq|\mathcal{C}|\leq\frac{n}{\delta}$, there exists a partition of $V$ into $V_{0},V_{1},\ldots,V_{L}$, of $\mathcal{C}$ into $\mathcal{C}_{0},\mathcal{C}_{1},\ldots,\mathcal{C}_{M}$ and a spanning sub-digraph $D_{c}^{\prime}$ of $D_{c}$ for each $c\in\mathcal{C}$ such that the following properties hold:
@@
Given a digraph collection $\mathcal{D}=\{D_{i}:i\in\mathcal{C}\}$ on $V$ and parameters $\epsilon>0$, $d\in[0,1)$ and $L_{0}\geq 1$, the reduced digraph collection $\mathcal{R}=\mathcal{R}(\epsilon,d,L_{0})$ of $\mathcal{D}$ is defined as follows. Apply Lemma 2.2 to $\mathcal{D}$ with parameters $\epsilon,\delta,d,L_{0}$ to obtain $\mathcal{D}^{\prime}$ and a partition $V_{0},V_{1},\ldots,V_{L}$ of $V$, and $\mathcal{C}_{0},\mathcal{C}_{1},\ldots,\mathcal{C}_{M}$ of $\mathcal{C}$, where $V_{0},\mathcal{C}_{0}$ are the exceptional sets and $V_{1},\ldots,V_{L}$ are the vertex clusters and $\mathcal{C}_{1},\ldots,\mathcal{C}_{M}$ are the color clusters. Then :=$\mathcal{R}:=\{R_{1},\ldots,R_{M}\}$ is a digraph collection of $M$ digraphs each on the same vertex set $[L]$, where, for $\{(h,i),j\}\in\binom{[L]}{2}\times[M]$, we have ${hi}\in R_{j}$ if and only if $\mathcal{D}_{hi,j}^{\prime}$ is $(\epsilon,d)$-regular.
@@
Now let $i\in[L]$ and $v\in V_{i}$. For each cluster $\mathcal{C}_{j}$ as above, choose an arbitrary color $c_{j}\in\mathcal{C}_{j}\cap\mathcal{D}_{v}$. Thus, the number of clusters $V_{h}$ containing some vertices in $N_{D_{c_{j}}^{\prime}}^{+}(v)$ is at least
@@
Lemma 2.2 (v) implies that $i$ is adjacent to each such $V_{h}$ in $R_{j}$. Therefore, for every $i\in[L]$, $d_{R_{j}}^{+}(i)\geq(p+\frac{\gamma}{2})L$ for at least $(1-\frac{d^{\frac{1}{4}}}{2})M$ colors $j$. The proofs of other cases are similar and we omit it. ∎
@@ 2.2 Transversal perfect matching in bipartite graph collections
Let $D$ be a digraph with vertex set $V$. We define $B_{D}$ to be the characteristic bipartite graph of $D$ with bipartition $V_{1}\cup V_{2}$ where $V_{1}=V_{2}=V$, for each $u\in V_{1}$ and $v\in V_{2}$, $uv\in E(B_{D})$ if and only if $uv\in E(D)$. Let $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ be a collection of digraphs on a common vertex set $V$. Define the characteristic bipartite graph collection of $\mathcal{D}$ to be $\mathcal{B_{D}}=\{B_{D_{1}},\ldots,B_{D_{n}}\}$.
A matching in a (di)graph $G$ is a collection of vertex-disjoint edges $M\subseteq E(G)$. We say that $M$ is a perfect matching if $V(M)=V(G)$. It is routine to check that $\mathcal{B_{D}}$ contains a transversal perfect matching if and only if the vertex set of $\mathcal{D}$ can be covered by a set of disjoint rainbow cycles inside $\mathcal{D}$. Therefore, in the following of this section, we are to find a transversal perfect matching in a bipartite graph collection.
@@
Let $\epsilon>0$ and $G$ be a balanced bipartite graph on vertex set $V_{1}\cup V_{2}$ of size $2n$. We say that
@@
    $A_{i},B_{i},C_{i}\subseteq V_{i}$, $|A_{i}|=|B_{i}|=(\frac{1}{2}-\epsilon)n$ and $|C_{i}|=2\epsilon n$ for $i\in[2]$,
@@
Let $10<\frac{1}{n},\epsilon,\delta<$ 1 where $n\in\mathbb{N}$ and let $\mathcal{G}=\{G_{1},\ldots,G_{n}\}$ be a balanced bipartite graph collection on a common vertex set $V$ of size $2n$. Given $i,j\in[n]$ such that $G_{i}$ and $G_{j}$ are both $\epsilon$-extremal, assume that the characteristic partition corresponding to $G_{\ell}$ is $(A_{1}^{\ell},B_{1}^{\ell},C_{1}^{\ell},A_{2}^{\ell},B_{2}^{\ell},C_{2}^{\ell})$ for $\ell\in\{i,j\}$. We say that they are $\delta$-crossing if $|A^{i}_{1}\Delta A^{j}_{1}|\geq\delta n$ and $|A^{i}_{1}\Delta B^{j}_{1}|\geq\delta n$.
@@
Let $10<\gamma,\alpha,\epsilon,\delta<$ 1. Suppose that $\mathcal{G}=\{G_{1},\ldots,G_{n}\}$ is a collection of bipartite graphs on a common vertex partition $V_{1}\cup V_{2}$ with $|V_{1}|=|V_{2}|=n$. We say that
@@ 3 Absorbing
Lemma 3.1 ([10]).
Suppose that $k,C,n\in\mathbb{N}$, $0\leq\frac{1}{n}\ll\gamma\ll\epsilon\ll\frac{1}{k},\frac{1}{C}\leq 1$, $m\in[n^{C}]$ and $t:=\gamma n$. Let $\mathbf{H}=\{H_{1},\ldots,H_{t}\}$ be a collection of directed $k$-graphs and let $\mathbf{Z}=\{Z_{1},\ldots,Z_{m}\}$ be a collection of directed multi-$k$-graphs all defined on a common vertex set $V$ of size $n$. Suppose that $|E(H_{i})|\geq\epsilon n^{k}$ for all $i\in[t]$, and for each $j\in[m]$ we have $|E(Z_{j})\cap E(H_{i})|\geq\epsilon n^{k}$ for at least $\epsilon t$ indices $i\in[t]$. Then there is a rainbow matching $M$ in $\mathbf{H}$ of size at least $(1-\frac{\epsilon^{2}}{4})t$ and $|E(Z_{j})\cap E(M)|\geq\frac{\epsilon^{2}t}{4}$ for each $j\in[m]$.
We now define the extremality for a single digraph.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
Definition 3.4 (directed absorbing path and absorbing cycle).
@@
If there is no ambiguity, we typically refer to the Type-I and Type-II directed $c$-absorbing path (cycle) simply as directed $c$-absorbing path (cycle).
@@
Let $10<\gamma,\alpha<$ 1. Suppose that $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ is a collection of digraphs on a common vertex set $V$ of size $n$. We say that $\mathcal{D}$ is $(\gamma,\alpha)$-strongly stable if $D_{i}$ is $\alpha$-nice for at least $\gamma n$ colors $i\in[n]$.
The following lemma finds a directed absorbing cycle when $\mathcal{D}$ is strongly stable.
@@
Let $1<\frac{1}{n}\ll\lambda,\mu\ll\gamma,\alpha\ll 1$. Assume that $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ is a collection of digraphs on a common vertex set $V$ of size $n$ with $\delta^{0}(\mathcal{D})\geq\frac{1-\mu}{2}n$. If $\mathcal{D}$ is $(\gamma,\alpha)$-strongly stable, then there exists a Type-I directed absorbing cycle with parameters $(1,0,\lambda,\lambda^{2})$.
@@
Since $\mathcal{D}$ is $(\gamma,\alpha)$-strongly stable, one may assume that $D_{1},\ldots,D_{\gamma n}$ are $\alpha$-nice and $6|\lambda n$. We divide the color set $[\frac{\lambda n}{2}]$ into consecutive sets $\{1,2,3\},\{4,5,6\},\ldots,\{\frac{\lambda n}{2}-2,\frac{\lambda n}{2}-1,\frac{\lambda n}{2}\}$. For each $i\in I:=[\frac{\lambda n}{6}]$, define $\mathcal{F}_{i}$ to be a directed $4$-graph with vertex $V$ and edge set
@@
By relabeling indices, one may assume $I^{\prime}=:[s]$ with $s=(1-\frac{\alpha^{2}}{100})\frac{\lambda n}{6}$. We are to connect the directed paths $P^{1},\ldots,P^{s}$ one by one into a rainbow directed cycle $C$. Firstly, we connect $P^{1}$ and $P^{2}$ into a single rainbow directed path $P^{1}xyP^{2}$. For this, choose distinct unused colors $c_{1},c_{2}\in[\lambda n+1,n]$ and $c_{3}\in[\frac{\lambda n}{2}+1,\lambda n]$. Let $U:=V\setminus V(\bigcup_{i\in[s]}P^{i})$. Clearly, $|U|\geq n-\lambda n$. It follows from $\delta^{0}(\mathcal{D})\geq(\frac{1-\mu}{2})n$ that
@@
We apply the above process for each pair $(P^{i},P^{i+1})$ for each $i\in[s]$ with $P^{s+1}:=P^{1}$. In each process, we need two unused colors from $[\lambda n+1,n]$ and one unused color from $[\frac{\lambda n}{2}+1,\lambda n]$. This is feasible since $2s\leq\frac{\lambda n}{3}$. The total number of used vertices at each step is at most $6s\leq{\lambda n}$, so we have the same degree bounds as (1). It is straightforward to verify that the resulting directed cycle $C$ is a Type-I directed absorbing cycle with parameters $(1,0,\frac{6s}{n},\frac{\alpha^{2}\lambda}{600})$ and hence with parameters $(1,0,\lambda,\lambda^{2})$ since $\lambda\ll\alpha$. ∎
The extremality of a digraph collection also depends on the relationships between the digraphs, so we give the following definitions.
@@
Let $10<\frac{1}{n},\epsilon,\delta<$ 1 with $n\in\mathbb{N}$. Let $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ be a collection of digraphs on a common vertex set $V$ of size $n$. Choose $i,j\in[n]$ such that $D_{i}$ is $(\epsilon,{\rm EC}t)$-extremal and $D_{j}$ is $(\epsilon,{\rm EC}s)$-extremal with $t,s\in[3]$. We say that $D_{i}$ and $D_{j}$ are $\delta$-crossing if one of the following holds:
@@
    $t,s\in\{1,2\}$, $|A_{i}\Delta A_{j}|\geq\delta n$ and $|A_{i}\Delta B_{j}|\geq\delta n$;
@@
    $t=s=3$, either $|W_{i}^{1}\Delta W_{j}^{1}|\geq\delta n$ and $|W_{i}^{1}\Delta W_{j}^{3}|\geq\delta n$, or $|W_{i}^{2}\Delta W_{j}^{2}|\geq\delta n$ and $|W_{i}^{2}\Delta W_{j}^{4}|\geq\delta n$;
@@
    $t\in\{1,2\},s=3$, either $|A_{i}\Delta W_{j}^{1}|\geq\delta n$ and $|A_{i}\Delta W_{j}^{3}|\geq\delta n$, or $|A_{i}\Delta W_{j}^{2}|\geq\delta n$ and $|A_{i}\Delta W_{j}^{4}|\geq\delta n$.
For $k\in[3]$, define the cross graph $C_{\mathcal{D},k}^{\epsilon,\delta}$ to be the graph with vertex set $[n]$ where $i$ and $j$ are adjacent if and only if $D_{i},D_{j}$ are satisfying (Ak).
We say that $\mathcal{D}$ is $(\epsilon,\delta)$-weakly stable if $e(C_{\mathcal{D},k}^{\epsilon,\delta})\geq\delta n^{2}$ for some $k\in[3]$.
@@
Let $10<\frac{1}{n}\ll\lambda,\mu,\epsilon\ll\delta<$ 1 and $\delta\ll\min\{\zeta,\frac{1}{2}-\zeta-\epsilon\}$ (where $\zeta$ is defined in Lemma 3.3). Assume that $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ is a collection of digraphs on a common vertex set $V$ of size $n$ with $\delta^{0}(\mathcal{D})\geq\frac{1-\mu}{2}n$. If $\mathcal{D}$ is $(\epsilon,\delta)$-weakly stable, then there exists a directed absorbing cycle with parameters $(\frac{\delta}{3},\sqrt{\epsilon},\lambda,\lambda^{2})$.
@@
Since $\mathcal{D}$ is $(\epsilon,\delta)$-weakly stable, one has $e(C_{\mathcal{D},k}^{\epsilon,\delta})\geq\delta n^{2}$ for some $k\in[3]$. In particular, if $k=2$, then we consider only $|W_{i}^{1}\Delta W_{j}^{1}|\geq\delta n$ and $|W_{i}^{1}\Delta W_{j}^{3}|\geq\delta n$ for all $ij\in E(C_{\mathcal{D},2}^{\epsilon,\delta})$; if $k=3$, then we consider only $|A_{i}\Delta W_{j}^{1}|\geq\delta n$ and $|A_{i}\Delta W_{j}^{3}|\geq\delta n$ for all $ij\in E(C_{\mathcal{D},3}^{\epsilon,\delta})$ (since the other case can be discussed similarly).
@@
We first give the following claim, which requires only the minimum semi-degree condition.
@@
$d_{i}^{-}(v,A_{j})\geq|A_{i}\cap A_{j}|-(|A_{i}|-d_{i}^{-}(v,A_{i}))\geq\frac{\delta n}{4}-\epsilon n\geq\frac{\delta n}{5}\ (\text{since}\ \epsilon\ll\delta).$
@@
[…]

### v2

@@ 2.1 Regularity for digraph collections
For every integer $L_{0}\geq 1$ and every $\epsilon,\delta>0$, there is an integer $n_{0}:=n_{0}(\epsilon,\delta,L_{0})$ such that for every $d\in[0,1)$ and every digraph collection $\mathcal{D}=\{D_{i}:i\in\mathcal{C}\}$ on vertex set $V$ of size $n\geq n_{0}$ with $\delta n\leq|\mathcal{C}|\leq\frac{n}{\delta}$, there exists a partition of $V$ into $V_{0},V_{1},\ldots,V_{L}$, of $\mathcal{C}$ into $\mathcal{C}_{0},\mathcal{C}_{1},\ldots,\mathcal{C}_{M}$ and a spanning sub-digraph $D_{c}^{\prime}$ of $D_{c}$ for each $c\in\mathcal{C}$ satisfying the following properties:
@@
Given a digraph collection $\mathcal{D}=\{D_{i}:i\in\mathcal{C}\}$ on $V$ and parameters $\epsilon>0$, $d\in[0,1)$ and $L_{0}\geq 1$, the reduced digraph collection $\mathcal{R}:=\mathcal{R}(\epsilon,d,L_{0})$ of $\mathcal{D}$ is defined as follows. Apply Lemma 2.2 to $\mathcal{D}$ with parameters $\epsilon,\delta,d,L_{0}$ to obtain $\mathcal{D}^{\prime}$, a partition $V_{0},V_{1},\ldots,V_{L}$ of $V$, and a partition $\mathcal{C}_{0},\mathcal{C}_{1},\ldots,\mathcal{C}_{M}$ of $\mathcal{C}$, where $V_{0},\mathcal{C}_{0}$ are the exceptional sets and $V_{1},\ldots,V_{L}$ are the vertex clusters and $\mathcal{C}_{1},\ldots,\mathcal{C}_{M}$ are the color clusters. Then $\mathcal{R}:=\{R_{1},\ldots,R_{M}\}$ is a digraph collection of $M$ digraphs each on the same vertex set $[L]$, where, for $\{(h,i),j\}\in\binom{[L]}{2}\times[M]$, we have ${hi}\in R_{j}$ if and only if $\mathcal{D}_{hi,j}^{\prime}$ is $(\epsilon,d)$-regular.
@@
Now let $i\in[L]$ and $v\in V_{i}$. For each cluster $\mathcal{C}_{j}$ as above, choose an arbitrary color $c_{j}\in\mathcal{C}_{j}\cap\mathcal{D}_{v}$. Thus, the number of vertex clusters $V_{h}$ containing some vertices in $N_{D_{c_{j}}^{\prime}}^{+}(v)$ is at least
@@
Lemma 2.2 (v) implies that $i$ is adjacent to each such $V_{h}$ in $R_{j}$. Therefore, for every $i\in[L]$, $d_{R_{j}}^{+}(i)\geq(p+\frac{\gamma}{2})L$ for at least $(1-\frac{d^{\frac{1}{4}}}{2})M$ colors $j$. The proofs of other cases are similar and we omit the proof. ∎
@@ 2.2 Transversal perfect matching in bipartite graph collections
In the study of digraphs, a natural approach is to consider their characteristic bipartite graphs. This subsection introduces the corresponding characteristic bipartite graph collection associated with a digraph system, and studies the existence of transversal perfect matchings in this bipartite system.
Let $D$ be a digraph with vertex set $V$. We define $B_{D}$ to be the characteristic bipartite graph of $D$ with bipartition $V_{1}\cup V_{2}$ where $V_{1}=V_{2}=V$; for each $u\in V_{1}$ and $v\in V_{2}$, $uv\in E(B_{D})$ if and only if $uv\in E(D)$. Let $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ be a collection of digraphs on a common vertex set $V$. Define the characteristic bipartite graph collection of $\mathcal{D}$ to be $\mathcal{B_{D}}=\{B_{D_{1}},\ldots,B_{D_{n}}\}$. A matching in a (di)graph $G$ is a collection of vertex-disjoint edges $M\subseteq E(G)$. We say that $M$ is a perfect matching if $V(M)=V(G)$.
It is straightforward to check that $\mathcal{B_{D}}$ contains a transversal perfect matching if and only if the vertex set of $\mathcal{D}$ can be covered by a set of disjoint rainbow cycles inside $\mathcal{D}$. Therefore, in the remainder of this section, we are to find a transversal perfect matching in a bipartite graph collection.
@@
Let $\epsilon>0$ and $G$ be a balanced bipartite graph on the vertex set $V_{1}\cup V_{2}$ of size $2n$. We say that
@@
    $A_{i},B_{i},C_{i}\subseteq V_{i}$, $|A_{i}|=|B_{i}|=(\frac{1}{2}-\epsilon)n$ and $|C_{i}|=2\epsilon n$ for each $i\in[2]$,
@@
Let $0<\frac{1}{n},\epsilon,\delta<1$ where $n\in\mathbb{N}$ and let $\mathcal{G}=\{G_{1},\ldots,G_{n}\}$ be a balanced bipartite graph collection on a common vertex set $V$ of size $2n$. Given $i,j\in[n]$ such that $G_{i}$ and $G_{j}$ are both $\epsilon$-extremal, assume that the characteristic partition corresponding to $G_{\ell}$ is $(A_{1}^{\ell},B_{1}^{\ell},C_{1}^{\ell},A_{2}^{\ell},B_{2}^{\ell},C_{2}^{\ell})$ for $\ell\in\{i,j\}$. We say that they are $\delta$-crossing if $|A^{i}_{1}\triangle A^{j}_{1}|\geq\delta n$ and $|A^{i}_{1}\triangle B^{j}_{1}|\geq\delta n$.
@@
Let $0<\gamma,\alpha,\epsilon,\delta<1$. Suppose that $\mathcal{G}=\{G_{1},\ldots,G_{n}\}$ is a collection of bipartite graphs on a common vertex partition $V_{1}\cup V_{2}$ with $|V_{1}|=|V_{2}|=n$. We say that
@@ 3 Absorbing
In this section, we first define the extremality for a single digraph and distinguish the characteristic partition of such digraphs. We then introduce our absorbing structure and prove that every digraph collection contains a directed absorbing cycle whenever either a large proportion of its members are non-extremal, or almost all of them are extremal while their characteristic partitions differ significantly.
@@
    there is a partition $A\cup B\cup L$ of $V$ such that $|A|=|B|=(\frac{1}{2}-\epsilon)n$, $d^{+}(a,{A}),d^{-}(a,{A})\geq(\frac{1}{2}-2\epsilon)n$ for all $a\in{A}$, $d^{+}(b,{B}),d^{-}(b,B)\geq(\frac{1}{2}-2\epsilon)n$ for all $b\in{B}$ and either $e(A,B)\leq\epsilon n^{2}$ or $e(B,A)\leq\epsilon n^{2}$; here we say $D$ is $(\epsilon,{\rm EC}1)$-extremal,
@@
    there is a partition $A\cup B\cup L$ of $V$ such that $|A|=|B|=(\frac{1}{2}-\epsilon)n$, $d^{+}(a,B),d^{-}(a,B)\geq(\frac{1}{2}-2\epsilon)n$ for all $a\in{A}$, $d^{+}(b,A),d^{-}(b,A)\geq(\frac{1}{2}-2\epsilon)n$ for all $b\in{B}$ and either $e(A)\leq\epsilon n^{2}$ or $e(B)\leq\epsilon n^{2}$; here we say $D$ is $(\epsilon,{\rm EC}2)$-extremal,
@@
    there is a constant $\zeta$ with $\epsilon^{1/3}\leq\zeta\leq\frac{1}{2}-\epsilon-\epsilon^{1/3}$ and a partition $C_{1}\cup C_{2}\cup C_{3}\cup C_{4}\cup L$ of $V$ such that $|C_{1}|=|C_{3}|=\zeta n$, $|C_{2}|=|C_{4}|=(\frac{1}{2}-\zeta-\epsilon)n$, for all $a\in C_{i}$ we have $d^{+}(a,C_{i})\geq(\zeta-\epsilon)n$ if $i\in\{1,3\}$, $d^{+}(a,C_{6-i})\geq(\frac{1}{2}-\zeta-2\epsilon)n$ if $i\in\{2,4\}$, and $d^{+}(a,C_{i+1})\geq(|C_{i+1}|-\epsilon)n$ for all $i\in[4]$ (here we identify $C_{5}$ with $C_{1}$), and either $e(C_{1},C_{3})\leq\epsilon n^{2}$ or $e(C_{3},C_{1})\leq\epsilon n^{2}$, either $e(C_{2})\leq\epsilon n^{2}$ or $e(C_{4})\leq\epsilon n^{2}$; here we say $D$ is $(\epsilon,{\rm EC}3)$-extremal, (see Figure 1 for the structures of EC1, EC2 and EC3).
@@
We call the partition obtained in Lemma 3.2 a characteristic partition of $D$. Assume that $0<\frac{1}{n}\ll\mu\ll\epsilon\leq 1$ and $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ is a collection of digraphs on a common vertex set $V$ of size $n$ with $\delta^{0}(\mathcal{D})\geq(\frac{1}{2}-\mu)n$. In view of Lemma 3.2, whenever $D_{i}$ is $\epsilon$-extremal, one may fix a characteristic partition $(A_{i},B_{i},L_{i})$ of $D_{i}$ if it is $(\epsilon,{\rm EC}k)$-extremal for some $k\in[2]$, a characteristic partition $(C_{i}^{1},C_{i}^{2},C_{i}^{3},C_{i}^{4},L_{i})$ of $D_{i}$ if it is $(\epsilon,{\rm EC}3)$-extremal. For convenience, define $W_{i}^{j}:=C_{i}^{j}\cup C_{i}^{j+1}$ for $j\in[4]$ (here we identify $C_{i}^{5}$ with $C_{i}^{1}$). We say that a vertex $v\in V$ is $D_{i}$-good if either $D_{i}$ is not $\epsilon$-extremal or $D_{i}$ is $\epsilon$-extremal and $v\in V\setminus L_{i}$. For every pair of distinct vertices $x,y\in V$, we define $L({xy}):=\{i\in[n]:{xy}\in E(D_{i})\}$ to be the set of colors appearing on xy{xy}.
Definition 3.3 (directed absorbing path and directed absorbing cycle).
@@
If there is no ambiguity, we typically refer to the Type-I and Type-II directed $c$-absorbing path (cycle) simply as a directed $c$-absorbing path (cycle).
[…]
@@
Lemma 3.4 ([9]).
Suppose that $k,C,n\in\mathbb{N}$, $0<\frac{1}{n}\ll\gamma\ll\epsilon\ll\frac{1}{k},\frac{1}{C}\leq 1$, $m\in[n^{C}]$ and $t:=\gamma n$. Let $\mathbf{H}=\{H_{1},\ldots,H_{t}\}$ be a collection of directed $k$-graphs and let $\mathbf{Z}=\{Z_{1},\ldots,Z_{m}\}$ be a collection of directed multi-$k$-graphs all defined on a common vertex set $V$ of size $n$. Suppose that $|E(H_{i})|\geq\epsilon n^{k}$ for all $i\in[t]$, and for each $j\in[m]$ we have $|E(Z_{j})\cap E(H_{i})|\geq\epsilon n^{k}$ for at least $\epsilon t$ indices $i\in[t]$. Then there is a rainbow matching $M$ in $\mathbf{H}$ of size at least $(1-\frac{\epsilon^{2}}{4})t$ and $|E(Z_{j})\cap E(M)|\geq\frac{\epsilon^{2}t}{4}$ for each $j\in[m]$.
We first consider the case where a digraph collection contains many non-extremal digraphs; such a collection is said to be strongly stable.
@@
Let $0<\gamma,\alpha<1$. Suppose that $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ is a collection of digraphs on a common vertex set $V$ of size $n$. We say that $\mathcal{D}$ is $(\gamma,\alpha)$-strongly stable if $D_{i}$ is $\alpha$-nice for at least $\gamma n$ colors $i\in[n]$.
By applying Lemma 3.4, we establish the following lemma, which yields a directed absorbing cycle when $\mathcal{D}$ is strongly stable.
@@
Let $0<\frac{1}{n}\ll\lambda,\mu\ll\gamma,\alpha\ll 1$. Assume that $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ is a collection of digraphs on a common vertex set $V$ of size $n$ with $\delta^{0}(\mathcal{D})\geq\frac{1-\mu}{2}n$. If $\mathcal{D}$ is $(\gamma,\alpha)$-strongly stable, then there exists a Type-I directed absorbing cycle with parameters $(1,0,\lambda,\lambda^{2})$.
@@
Since $\mathcal{D}$ is $(\gamma,\alpha)$-strongly stable, one may assume that $D_{1},\ldots,D_{\gamma n}$ are $\alpha$-nice and $6|\lambda n$. We divide the color set $\frac{\lambda n}{2}]$ into consecutive sets $\{1,2,3\},\{4,5,6\},\ldots,\{\frac{\lambda n}{2}-2,\frac{\lambda n}{2}-1,\frac{\lambda n}{2}\}$. For each $i\in I:=[\frac{\lambda n}{6}]$, define $\mathcal{F}_{i}$ to be a directed $4$-graph with vertex $V$ and edge set
@@
By relabeling indices, one may assume $I^{\prime}=:[s]$ with $s=(1-\frac{\alpha^{2}}{100})\frac{\lambda n}{6}$. We are to connect the directed paths $P^{1},\ldots,P^{s}$ one by one into a rainbow directed cycle $C$. Firstly, we connect $P^{1}$ and $P^{2}$ into a single rainbow directed path $P^{1}xyP^{2}$. For this, choose distinct unused colors $c_{1},c_{2}\in[\lambda n+1,n]$ and $c_{3}\in[\frac{\lambda n}{2}+1,\lambda n]$. Let $U:=V\setminus V(\bigcup_{i\in[s]}P^{i})$. Clearly, $|U|\geq n-\lambda n$. It follows from $\delta^{0}(\mathcal{D})\geq\frac{1-\mu}{2}n$ that
@@
We apply the above process for each pair $(P^{i},P^{i+1})$ for each $i\in[s]$ with $P^{s+1}:=P^{1}$. In each process, we need two unused colors from $\lambda n+1,n]$ and one unused color from $\frac{\lambda n}{2}+1,\lambda n]$. This is feasible since $2s\leq\frac{\lambda n}{3}$. The total number of used vertices at each step is at most $6s\leq{\lambda n}$, so we have the same degree bounds as (1). It is straightforward to verify that the resulting directed cycle $C$ is a Type-I directed absorbing cycle with parameters $(1,0,\frac{6s}{n},\frac{\alpha^{2}\lambda}{600})$ and hence with parameters $(1,0,\lambda,\lambda^{2})$ since $\lambda\ll\alpha$. ∎
The extremality of a digraph collection also depends crucially on the relationships between different digraphs. This leads us to introduce the following definitions.
@@
Let $0<\frac{1}{n}\ll\epsilon,\delta<1$ with $n\in\mathbb{N}$. Let $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ be a collection of digraphs on a common vertex set $V$ of size $n$. Choose $i,j\in[n]$ such that $D_{i}$ is $(\epsilon,{\rm EC}t)$-extremal and $D_{j}$ is $(\epsilon,{\rm EC}s)$-extremal with $t,s\in[3]$. We say that $D_{i}$ and $D_{j}$ are $\delta$-crossing if one of the following holds:
@@
    $t,s\in\{1,2\}$, $|A_{i}\triangle A_{j}|\geq\delta n$ and $|A_{i}\triangle B_{j}|\geq\delta n$;
@@
    $t=s=3$, either $|W_{i}^{1}\triangle W_{j}^{1}|\geq\delta n$ and $|W_{i}^{1}\triangle W_{j}^{3}|\geq\delta n$, or $|W_{i}^{2}\triangle W_{j}^{2}|\geq\delta n$ and $|W_{i}^{2}\triangle W_{j}^{4}|\geq\delta n$;
@@
    $t\in\{1,2\},s=3$, either $|A_{i}\triangle W_{j}^{1}|\geq\delta n$ and $|A_{i}\triangle W_{j}^{3}|\geq\delta n$, or $|A_{i}\triangle W_{j}^{2}|\geq\delta n$ and $|A_{i}\triangle W_{j}^{4}|\geq\delta n$.
For $k\in[3]$, define the cross graph $C_{\mathcal{D},k}^{\epsilon,\delta}$ to be the graph with vertex set $[n]$ where $i$ and $j$ are adjacent if and only if $D_{i},D_{j}$ are satisfying (Ak). We say that $\mathcal{D}$ is $(\epsilon,\delta)$-weakly stable if $e(C_{\mathcal{D},k}^{\epsilon,\delta})\geq\delta n^{2}$ for some $k\in[3]$.
The subsequent observation is an immediate consequence of the above definition.
@@
Let $0<\frac{1}{n}\ll\lambda,\mu\ll 8\epsilon^{1/2}\leq\delta\leq\epsilon^{1/3}\ll 1$. Assume that $\mathcal{D}=\{D_{1},\ldots,D_{n}\}$ is a collection of digraphs on a common vertex set $V$ of size $n$ with $\delta^{0}(\mathcal{D})\geq\frac{1-\mu}{2}n$. If $\mathcal{D}$ is $(\epsilon,\delta)$-weakly stable, then there exists a directed absorbing cycle with parameters $(\frac{\delta}{3},\sqrt{\epsilon},\lambda,\lambda^{2})$.
@@
Since $\mathcal{D}$ is $(\epsilon,\delta)$-weakly stable, one has $e(C_{\mathcal{D},k}^{\epsilon,\delta})\geq\delta n^{2}$ for some $k\in[3]$. In particular, if $k=2$, then we consider only $|W_{i}^{1}\triangle W_{j}^{1}|\geq\delta n$ and $|W_{i}^{1}\triangle W_{j}^{3}|\geq\delta n$ for all $ij\in E(C_{\mathcal{D},2}^{\epsilon,\delta})$; if $k=3$, then we consider only $|A_{i}\triangle W_{j}^{1}|\geq\delta n$ and $|A_{i}\triangle W_{j}^{3}|\geq\delta n$ for all $ij\in E(C_{\mathcal{D},3}^{\epsilon,\delta})$ (since the other case can be discussed similarly).
@@
In order to apply Lemma 3.4, we first give the following claim, which requires only the minimum semi-degree condition.
@@
$d_{i}^{-}(v,A_{j})\geq|A_{i}\cap A_{j}|-(|A_{i}|-d_{i}^{-}(v,A_{i}))\geq\frac{\delta n}{4}-\epsilon n\geq\frac{\delta n}{5}\ (\text{since}\ \epsilon\leq\delta^{2}).$
@@
Since $u$ is $D_{j}$-good, one has $u\in Y_{j}$ for some $Y\in\{A,B\}$. Let $Z=Y$ if $D_{j}$ is $(\epsilon,{\rm EC}1)$-extremal, and let $Z\in\{A,B\}\setminus\{Y\}$ if $D_{j}$ is $(\epsilon,{\rm EC}2)$-extremal. Notice that $|N_{j}^{+}(w,Z_{j})|\geq(\frac{1}{2}-2\epsilon)n$ for each vertex $w\in Y_{j}$. Choose $x\in N_{i}^{-}(v,Y_{j})$. Then